Problem:
Working together, pipe I and Pipe II can fill a tank in 12 hours. Pipe I takes twice as long as pipe II to fill the tank. How long would it take for pipe II, working by itself, to fill the tank.
solution: Suppose pipe II takes x hours,
then pipe I takes 2X hours.
together they take 12 hours.
so x + 2x = 12hours
3x = 12
x = 4
But I was told that my problem is wrong... and it is to solved using the formulae: work done = rate * time.
I got this problem in my exam so i do not have the ISBN.
Plz help me find answer for this
Thanking u in advance..

Answer by stanbon(75887)

Working together, pipe I and Pipe II can fill a tank in 12 hours. Pipe I takes twice as long as pipe II to fill the tank. How long would it take for pipe II, working by itself, to fill the tank.
-------------
You do the problem with rates, not with hours.
--------------
Pipes together DATA:
Time = 12 hr./job ; Rate = 1/12 job/hr
--------------------
Pipe II DATA:
Time = x hrs./job ; Rate = 1/x job/hr
----------------------
Pipe I DATA:
Time = 2x hrs./jog ; Rate = 1/2x job/hr
-----------------
EQUATION:
rate + rate = together rate
1/x + 1/2x = 1/12
Multiply thru by 12x to get:
12 + 6 = x
x=18 hrs (time it would take Pipe II to do the job)
================
Cheers,
Stan H.